The perimeter of a rectangle with width 10 inches and length 13 inches is
Answer by farohw(175)



The perimeter of a rectangle uses the formula: P = 2L + 2W

Therefore, P = 2(13) + 2(10) --> P = 26 + 20 = 46

The Perimeter (P) is 46 inches.